OpenStudy (anonymous):

H(x)= {((sin(5pix)-1)/(2x-1)) x=not1/2, k x=1/2}

OpenStudy (perl):

How can I help you :)

OpenStudy (anonymous):

\[h(x)=\left[\begin{matrix}\frac{ \sin(5 \pi x) - 1 }{ 2x-1 } & x \neq \frac{ 1 }{ 2 } \\ k & x \neq \frac{ 1 }{ 2 }\end{matrix}\right]\]

OpenStudy (anonymous):

Find the value of k that would make the function continuous in each case.

OpenStudy (perl):

ok that should say k , when x = 1/2

OpenStudy (perl):

we need to find limit [sin (5pix) - 1 ] / ( 2x-1) as x->1/2